PLEASE HELP ME WITH THIS "EXPRESS THE FOLLOWING REAL NUMBER AS THE QUOTIENT OF TWO INTEGERS"
1. 5 7/8
2. 25
3. 0
4. 10.62

Answer by stanbon(75887)

"EXPRESS THE FOLLOWING REAL NUMBER AS THE QUOTIENT OF TWO INTEGERS"
1. 5 7/8 = (8*5+7)/8 = 47/8
2. 25 = 25/1
3. 0 = 0/1
4. 10.62 = 1062/100
==========================
Cheers,
Stan H.